# Precipitation Downscaling with Spatiotemporal Video Diffusion

### v2

@@ 4 Related Work
Diffusion models [51, 17, 59, 53, 54] are a class of generative models which generate data via an iterative denoising process. Most closely related to our work are diffusion models for video. Recent models [63] propose to generate deterministic next-frame predictions autoregressively with additional residuals generated by a diffusion model, or generate videos directly via diffusion [15].
Image Super-Resolution
Single image super-resolution is a challenging and important task in computer vision [61]. While there exist many classical approaches to this problem [4, 10], deep learning based methods have recently become the dominant paradigm [61]. Supervised approaches are typical, where models are trained to minimize the error between generated and ground-truth images [65, 8, 30]. On the other hand, generative approaches [36, 46, 3] model the data distribution itself, often leading to higher quality outputs. In particular, cascaded diffusion models [39, 45] and SR3 [46] propose models for single image super-resolution but are unable to leverage the temporal context.
Video Super-Resolution
Many approaches for video super-resolution not based on diffusion models [6, 11, 19]. We refer to [33] for a more comprehensive overview. The recent VRT model [29] is a transformer-based model for video super-resolution with a recurrent variant RVRT [32] that focuses on parallel decoding and guided clip alignment. We note that these state-of-the-art approaches are deterministic, whereas our approach is generative, allowing us to directly model the (conditional) data distribution, thus preventing mode averaging.
@@
Advancements in climate data processing and downscaling have prominently featured the application of CNNs. [58] draw inspiration from FRVSR [47], adopting an iterative approach that utilizes the high-resolution frame estimated in the previous step as input for subsequent iterations. [62] employed Fourier neural operators for versatile downscaling. [13] contribute by generating physically consistent downscaled climate states, employing a softmax layer to enforce conservation laws.
A concurrent work [35] employs diffusion models for downscaling of climate states at each time point. The use of GANs has also been pervasive in downscaling and precipitation prediction, as extensively reviewed in recent literature [27, 38, 14, 40, 12, 60]. Beyond downscaling, [1] demonstrate the diffusion model’s efficacy in synthesizing full rain density from vorticity inputs. Additionally, [16] uses a diffusion model for downscaling solar irradiance. These models are also used for probabilistic weather forecasting and nowcasting [26, 28].
@@ 5 Conclusion
We propose a video super-resolution method for probabilistic precipitation downscaling. Our model, STVD, operates by deterministically super-resolving a given low-resolution frame sequence, followed by stochastically modeling the residual details via diffusion. Our model successfully resolves how fine-grid precipitation features, generated as weather systems, interact with complex coastlines and mountains based on temporally coherent coarse-grid information. Our method outperforms several competitive baselines on a range of quantitative metrics. This is an important step towards designing effective statistical downscaling methods to provide highly localized information about planning for extreme weather events, such as floods or hurricanes in a warming climate, based on computationally tractable coarse-grid atmospheric models.
@@ A.1 Model Architecture
Our architecture is a conditional extension of the DDPM [17] and SR3 [46] models. As discussed in Sec. 2, Fig. 5 outlines the architecture of the proposed denoising and downscaling networks. Furthermore, Fig. 6 outlines the detail of each box. Before elaborating on the details, we define the naming convention for the parameter choices that we adopt in this section:
@@
  • ResBlock: leverages a standard implementation of ResBlock, i.e., two blocks each with a weight-standardised convolution of $3\times 3$ kernel, Group Normalization over groups of 8 and SiLU activation, followed by a channel-adjusting $1\times 1$ convolution.
  • Attention: leverages standard implementation of quadratic or linear attention, incorporating 4 attention heads, each with a 32-dimensional representation. Utilizing a $3\times 3$ convolutional kernel, query, key, and value feature maps are generated, resulting in feature maps of size [b $\times$ t, h $\times$ c, x, y], where b, t, c, x, and y denote batch, time, channel, height, and width, respectively, with h representing the number of heads. These feature maps undergo rearrangement to the form [batch, head, channel, token]), facilitating self-attention between tokens of channel dimensions. The chosen rearrangement order determines whether spatial or temporal self-attention is performed. Subsequently, the feature maps revert to their original format, upon which a separate convolution operation is conducted, followed by layer normalization to project the feature maps back to their original dimensions, akin to their state before the initial convolution within the attention block. Further elaboration on the rearrangement choices for the variants is discussed, adhering to einops [^2] https://einops.rocks/1-einops-basics/ notation:
    • Q-Spatial: quadratic variant, applied in the bottleneck layer, self attends between every pixel of the feature map with the following rearrangement : [b $\times$ t, h $\times$ c, x, y] -> [b $\times$ t, h, c, x $\times$ y]
    • Q-Temporal: quadratic variant, applied in the bottleneck layer, self attends between feature maps across time with the following rearrangement : [b $\times$ t, h $\times$ c, x, y] -> [b, h, c $\times$ x $\times$ y, t]
    • L-Spatial: linear variant, applied in expansive and contractive layers of UNet, self attends between every pixel within a patch of the feature map with the following rearrangement : [b $\times$ t, h $\times$ c, x $\times$ p, y $\times$ p] -> [b $\times$ t, h $\times$ x $\times$ y, c, p $\times$ p] where p is patch size, starting with 192, halving at each contractive layer and doubling at each expansive layer.
    • L-Temporal: linear variant, applied in expansive and contractive layers of UNet, self attends between feature maps across time in a channel factorised manner with the following rearrangement : [b $\times$ t, h $\times$ c, x, y] -> [b, h $\times$ x $\times$ y, c, t]
  • MLP: conditioning on the denoising step $n$ is achieved through this block which uses 32-dimensional random Fourier features, followed a linear layer, GELU activation and another linear layer to transform the noise step to a higher dimension.
@@
Figure 5: The figure depicts the overall model architecture where the top UNet performs diffusion on the residual, conditioned on the noise step $n$, $\mathbf{x}^{0:T}$, $\mathbf{\bar{y}}^{0:T}$ and the context of the bottom U-Net feature map (refer to Eq. 3 in Sec. 2.1). The bottom UNet is the determinsitic downscaler (Eq. 1). Details of each block are shown in Fig. 6.
Figure 6: The details of the components of the modules shown in Fig. 5; the colored arrows in the modules correspond to the arrows with the same color in Fig. 5.
Fig. 5 shows how the U-Net from both the denoising and downscaling networks interact with each other. The top U-Net depicts the denoising network whereas the bottom U-Net depicts the downscaling network. The denoising network is conditioned in three ways. First, the network is conditioned on the bicubically downscaled low-resolution frame $\mathbf{x}^{0:T}$ and concatenating it with both $\mathbf{r}^{0:T}_{n}$ (noisy residual) and $\mathbf{\bar{y}}^{0:T}$ (downscaler output) along the channel dimension. Second, the network is also conditioned on the feature maps generated by the downscaler network (see Eq. 1 and Eq. 3) as shown via the green arrows connecting the downsampling units of both the networks. As detailed in the figure, L-(Spatial/Temporal) Attention blocks from contractive layers of the downscaler network yield a feature map that gets concatenated along the inputs of both ResBlocks of the contractive layer of the denoising network. Finally, each contractive and expansive layer of the diffusion UNet gets conditioned on the denoising step $n$ as shown via the black arrows. This conditional embedding for the step is generated through MLP. As shown in the Figure, this embedding is received by both ResBlocks. The information flows from the noisy residual through the network, as shown via blue arrows to predict the angular parameter $\mathbf{v}$. Both U-Nets have skip connections (shown via red arrows) between both ResBlocks of contractive and expansive layers of the same UNet.
@@
Table 2: FV3GFS uses a cubed-sphere grid, in which the surface of the globe is divided into six tiles. A single high-resolution tile covers 25km and is $384\times 384$. Our UNet Encoder and Decoder has 6 layers with a base channel dimension of 64 and multipliers as stated above.
Figure 7: An illustration of the training and inference pipelines of Swin-IR-Diff. Similar to Fig. 1, blue blocks represent operations common to both training and inference phases. Red blocks signify operations exclusive to training, while green blocks indicate inference-only processes. However, in a major contrast, it is an image-only downscaler. In summary, this model takes in a current low-resolution frame which is deterministically downscaled via Swin-IR, followed by modeling of the residual details via conditional diffusion. The model details remain similar to what is described in Appendix A.1 with the absence of downscaler-conditioning and temporal attention in the diffusion model.
@@ A.2 On Swin-IR-Diff and Multiple Channels
Here, we discuss SwinIR-Diff, which expands on one of our robust baselines. Sec. 3.3 provides a concise overview of Swin-IR-Diff. Shown as a sketch in Fig. 7, this model opts to downscale each precipitation state individually, akin to an image super-resolution model. Resembling SR3 in its foundation of a conditional diffusion model, Swin-IR-Diff adopts a residual pipeline. It involves a deterministic prediction corrected by a residual generated from the conditional diffusion model, with the Swin-IR model serving as the deterministic downscaler in this context.
@@ A.3 Additional Samples
In addition to illustrating precipitation downscaling in the Sierras and Central California, we present our model’s output for another unique region—the Himalayas. Fig. 8 mirrors Fig. 2, displaying outputs from different models. Note that for the same regional topography, Fig. 4 (left) compares the annual precipitation time average.
@@ A.3.1 Spectra
In Fig. 9, we plot (in log-scale) the squared-magnitude of the complex-valued FFT applied to an image in our evaluation set. Overall, we see that the samples from STVD closely match the ground-truth high resolution spectrum. The baselines RVRT, VSRT and PSRT demonstrate a spectrum which decays too rapidly, placing too little energy in the high-frequency components. Additionally, we see a banding in these spectrums similar to the one observed in ground-truth low resolution spectrum. This indicates that these baselines are overly smooth compared to the ground-truth and follow similar banding patterns of the corresponding low resolution image. For the Swin-IR baseline, we observed outliers of large magnitudes in the generated precipitation maps, which we hypothesize leads to the observed checkerboard pattern seen in the spectrum. Swin-IR-Diff seems to decay the spectra more rapidly than STVD.

### v3

@@ 4 Related Work
Diffusion models [56, 18, 59, 39, 40, 36] are a class of generative models based on an iterative denoising process. Closely related to our work are diffusion models for video. Recent models [68] generate deterministic next-frame predictions autoregressively with additional residuals generated by a diffusion model, or generate videos directly in pixel space [17, 64, 19] or in a latent space [5, 4]. While some works on video diffusion [5, 19] employ video super-resolution as a step in the overall modeling process, our work focuses exclusively on the video super-resolution task, particularly within the context of precipitation downscaling.
Super-Resolution
Within the computer vision community, the paradigm for single image super-resolution has shifted from classical approaches [3, 12] to deep learning based methods [66]. Generative approaches, like cascaded diffusion [44, 50], SR3 [51], and DiffPIR [73] employ diffusion models for image super-resolution. However, these are unable to leverage temporal context. On the other hand, many approaches for video super-resolution have been proposed [8, 13, 20]. For a more comprehensive overview, see [34]. Recent models of note include the transformer-based models PSRT [55] and VRT [30], as well as the recurrent variant RVRT [33], which focuses on parallel decoding and guided clip alignment. We emphasize that these state-of-the-art approaches are deterministic, where our approach is generative. This allows us to preventing mode averaging and to produce more realistic samples, which is particularly critical in the context of precipitation modeling.
@@
Recent years have seen advancements in data-driven climate and weather modeling [46, 38], with models like GraphCast [27], GenCast [43], and FourCastNet [41] providing forecasts that are competitive with meteorological methods while being significantly faster. Rather than replacing numerical forecasting methods, our approach seeks to augment their capabilities by downscaling coarse-grid predictions.
While downscaling for climate and weather has been approached using techniques based on domain knowledge [23, 22], we focus here on data-driven approaches. [63] draw inspiration from FRVSR [52], adopting an iterative approach that uses the high-resolution frame estimated in the previous step as input for subsequent iterations. [67] employed Fourier neural operators for downscaling at arbitrary resolutions. [15] generate physically consistent downscaled climate states, using a softmax layer to enforce conservation laws. These approaches, though, are deterministic and trained by minimizing the MSE, thus lacking the realism and uncertainty quantification provided by a generative approach.
In terms of generative approaches, concurrent work [37] employs diffusion models for downscaling climate states. The use of GANs has also been pervasive in downscaling and precipitation prediction [29, 42, 16, 45, 14, 65]. However, these GAN-based approaches inherit the mode collapse and training difficulties present in all GAN-based models [2]. Here, we highlight that these approaches are applied at each frame and cannot incorporate temporal information as is done in our model.
@@ 5 Conclusion
We propose a video super-resolution method for probabilistic precipitation downscaling. Our model, STVD, deterministically super-resolves a given low-resolution frame sequence and then stochastically models the residual details via diffusion. Our model effectively resolves how fine-grid precipitation features, generated as weather systems, interact with complex topography based on temporally coherent coarse-grid information. Our method outperforms several competitive baselines on a range of quantitative metrics. This is an important step towards designing effective statistical downscaling methods, providing highly localized information for planning against extreme weather events, such as floods or hurricanes in a warming climate, using tractable coarse-grid atmospheric models.
Limitations and Broader Impacts
A limitation of our approach is the necessity of paired low-resolution and high-resolution images for training. While this can be done once prior to training, designing methods that only require the low-resolution states is an interesting challenge. In terms of broader impacts, our approach could potentially have harmful consequences if adopted blindly to a new dataset, where distribution shift could cause the model performance to degrade, potentially leading to underestimation of extreme weather risks such as droughts or floods. To mitigate these risks, the model should be re-trained and rigorously evaluated on the dataset of interest.
@@ A.1 Model Architecture
Our architecture is a conditional extension of the DDPM [18] and SR3 [51] models. As discussed in Sec. 2, Fig. 6 presents the architecture of the proposed denoising and downscaling networks, while Fig. 7 provides detailed description of each component. Before elaborating on the specifics, we define the naming conventions for the parameter choices adopted in this section:
@@
  • ResBlock: a standard ResBlock implementation consisting of two blocks, each with a weight-standardised convolution using a $3\times 3$ kernel, Group Normalization over groups of 8 and SiLU activation, followed by a channel-adjusting $1\times 1$ convolution.
  • Attention: a standard implementation of quadratic or linear attention, incorporating 4 attention heads, each with a 32-dimensional representation. Using a $3\times 3$ convolutional kernel, query, key, and value feature maps are generated, resulting in feature maps of size [b $\times$ t, h $\times$ c, x, y], where b, t, c, x, and y denote batch, time, channel, height, and width, respectively, with h representing the number of heads. These feature maps are rearranged to [batch, head, channel, token]), facilitating self-attention between tokens of channel dimensions. The rearrangement order determines whether spatial or temporal self-attention is performed. Subsequently, the feature maps revert to their original format, upon which a separate convolution operation is conducted, followed by layer normalization to project the feature maps back to their original dimensions, akin to their state before the initial convolution within the attention block. Further elaboration on the rearrangement choices for the variants is discussed, adhering to einops [^2] https://einops.rocks/1-einops-basics/ notation:
    • Q-Spatial: quadratic variant applied in the bottleneck layer; self attends between every pixel of the feature map with the following rearrangement : [b $\times$ t, h $\times$ c, x, y] -> [b $\times$ t, h, c, x $\times$ y]
    • Q-Temporal: quadratic variant applied in the bottleneck layer; self attends between feature maps across time with the following rearrangement : [b $\times$ t, h $\times$ c, x, y] -> [b, h, c $\times$ x $\times$ y, t]
    • L-Spatial: linear variant applied in expansive and contractive layers of UNet; self attends between every pixel within a patch of the feature map with the following rearrangement : [b $\times$ t, h $\times$ c, x $\times$ p, y $\times$ p] -> [b $\times$ t, h $\times$ x $\times$ y, c, p $\times$ p] where p is patch size, starting with 192, halving at each contractive layer and doubling at each expansive layer.
    • L-Temporal: linear variant applied in expansive and contractive layers of UNet; self attends between feature maps across time in a channel factorised manner with the following rearrangement : [b $\times$ t, h $\times$ c, x, y] -> [b, h $\times$ x $\times$ y, c, t]
  • MLP: conditioning on the denoising step $n$ is achieved through this block, which uses 32-dimensional random Fourier features, followed by a linear layer, GELU activation and another linear layer to transform the noise step to a higher dimension.
@@
Figure 6: The figure depicts the overall model architecture where the top UNet performs diffusion on the residual, conditioned on the noise step $n$, $\mathbf{x}^{0:T}$, $\mathbf{\bar{y}}^{0:T}$ and the context of the bottom U-Net feature map (refer to Eq. 1 in Sec. 2.1). The bottom UNet is the deterministic downscaler. Details of each block are shown in Figure 7.
Figure 7: The details of the components of the modules shown in Figure 6; the colored arrows in the modules correspond to the arrows with the same color in Figure 6.
Fig. 6 illustrates the interaction between the U-Net architecture of the denoising and downscaling networks. The top U-Net depicts the denoising network, while the bottom U-Net depicts the downscaling network. The denoising network is conditioned in three ways. First, the network is conditioned on the bicubically downscaled low-resolution frames $\mathbf{x}^{0:T}$, concatenated with both the noisy residual $\mathbf{r}^{0:T}_{n}$ the downscaler output and $\mathbf{\bar{y}}^{0:T}$ along the channel dimension. Second, the network is also conditioned on the feature maps generated by the downscaler network as indicated by the green arrows connecting the downsampling units of both the networks. The L-(Spatial/Temporal) Attention blocks from contractive layers of the downscaler network yield a feature map that gets concatenated with the inputs of both ResBlocks of the contractive layer of the denoising network. Finally, each contractive and expansive layer of the diffusion UNet gets conditioned on the denoising step $n$, shown via the black arrows. This conditional embedding for the step is generated through MLP and received by both ResBlocks. The information flows from the noisy residual through the network as shown by the blue arrows to predict the angular parameter $\mathbf{v}$. Both U-Nets have skip connections, indicated by the red arrows, between both ResBlocks of contractive and expansive layers of the same UNet.
Table 2: FV3GFS uses a cubed-sphere grid, in which the surface of the globe is divided into six tiles. Each high-resolution tile covers 25 km and is $384\times 384$. Our UNet Encoder and Decoder have 6 layers with a base channel dimension of 64 and multipliers as stated above.
@@
Figure 8: An illustration of the training and inference pipelines of Swin-IR-Diff. Similar to Fig. 1, blue blocks represent operations common to both training and inference phases. Red blocks signify operations exclusive to training, while green blocks indicate inference-only processes. However, in contrast, it is an image-only downscaler. This model takes in a current low-resolution frame which is deterministically downscaled via Swin-IR, followed by modeling of the residual details via conditional diffusion. The model details remain similar to what is described in A.1 with the absence of downscaler-conditioning and temporal attention in the diffusion model.
Figure 9: Qualitative comparison between our proposed model and all baselines for a specific precipitation event in the Sierra mountain range. This figure is a repetition of Fig. 2 for better visual overview. The first row represents the ground truth fine-grid precipitation state sequence, and the last row represents the coarse-grid precipitation that is being downscaled. All other rows correspond to our model and the baseline outputs. The time interval between adjacent frames is 3 hours; the plotted region is $1000\times 1000$ km.
@@ A.2 On Swin-IR-Diff and Multiple Channels
Here, we discuss SwinIR-Diff, which expands on one of our robust baselines. Sec. 3 provides a concise overview of Swin-IR-Diff. Shown as a sketch in Fig. 8, this model opts to downscale each precipitation state individually, akin to an image super-resolution model. Resembling SR3 in its foundation of a conditional diffusion model, Swin-IR-Diff adopts a residual pipeline. It involves a deterministic prediction corrected by a residual generated from the conditional diffusion model, with the Swin-IR model serving as the deterministic downscaler in this context.
@@ A.3 Additional Samples
In addition to re-illustrating precipitation downscaling in the Sierras and Central California from Fig. 2 in Fig. 9), we present our model’s output for another unique region—the Himalayas. Fig. 10 mirrors Fig. 9, displaying outputs from different models. Note that for the same regional topography, Fig. 5 (left) compares the annual precipitation time average.
@@ A.3.1 Spectra
In Figure 11, we plot (in log-scale) the squared-magnitude of the complex-valued FFT applied to an image in our evaluation set. Overall, we see that the samples from STVD closely match the ground-truth high resolution spectrum. The baselines RVRT, VSRT and PSRT demonstrate a spectrum which decays too rapidly, placing too little energy in the high-frequency components. Additionally, we see a banding in these spectra. This indicates that these baselines are overly smooth compared to the ground-truth and follow similar banding patterns of the corresponding low resolution image. For the Swin-IR baseline, we observed outliers of large magnitudes in the generated precipitation maps, which we hypothesize leads to the observed checkerboard pattern seen in the spectrum. Swin-IR-Diff and VDM seem to decay the spectra more rapidly than STVD.